OpenStudy (anonymous):

It took a crew 1 h 48 min to row 6 km upstream and back again. If the rate of flow of the stream was 8 km/h, what was the rowing speed of the crew in still water?

OpenStudy (dumbcow):

d = r*t upstream: 6 = (v-8)t --> t = 6/(v-8) downstream: **Note 1 hr 48 min = 1.8 hr 6 = (v+8)(1.8-t) Substitute 6 = (v+8)(1.8 - 6/(v-8)) divide by v+8 6/v+8 = 1.8 - 6/v-8 6/v+8 + 6/v-8 = 1.8 combine fractions 6(v-8) +6(v+8) / (v^2-64) = 1.8 6(v-8) +6(v+8) = 1.8(v^2 -64) 12v = 1.8v^2 - 115.2 1.8v^2 -12v -115.2 = 0 use quadratic formula v = -5.33 or 12 speed must be positive v = 12